OpenStudy (anonymous):

Find the equation for the parabola with focus at (3, -3) and vertex at (3, 1).

OpenStudy (anonymous):

\[y-1= -4p (x-3)^2\\ p=1-(-3)=4 \]

OpenStudy (anonymous):

\[y = 1 -16(x-3)^2\\ y=-16 x^2+96 x-143 \]